Z. Given the following data representing the goods market in an open economy Marginal propensity to consume 0.5 300 Direct tax rate Autonomous Taxation Transfers 0.1 100 200 10000 Gov spending 2000 0.2 50 6000 investment Marrinal propensity to import Autonomous imports Exports Using the Keynesian cross-model, compute a) The equilibrium level of the aggregate output The value of public savings corresponding to the equilibrium level of the output; say if the country is running a budget deficit or surplus The value of net exports corresponding to the equilibrium level of the output; say if the country is running a trade deficit or surplus; c) d) The value of the aggregate output for which the system would run a trade balance; e) The variation in the equilibrium level of the aggregate output due to a variation in the government spending equal to Gs-1000 The new equilibrium level of the aggregate output if the marginal propensity to import decreases to 0.1
Answer #1

C = 300 + 0.5(Y - 0.1Y- 100 + 200 )

G =. 10,000

I = 2000

IMPORTS = 50 + 0.2Y

EXPORTS = 6000

a) Y = C + I + G + (X- M)

Y = 300 + 0.5(Y - 0.1Y- 100 + 200 ) + 2000 + 10,000 + 6000 - 50 -  0.2Y

Y= 18,250 + 0.5Y - 0.05Y - 50 + 100 - 0.2Y

Y= 18,300 - .25Y

.75Y = 18,300

Y = 24,400

B ) G = 10,000

transfers = 200

TAXES = 0.1Y + 100 = 0.1(24,400 ) + 100 = 2440 + 100 = 2540

since receipts (taxes )< expenditure ( government spending + transfers ) , government is running. a budget deficit

deficit = G + transfers - taxes

= 10,000 + 200 - 2540

= 7660

c) net exports = exports - imports

imports = 50 + 0.2Y = 50 + 0.2*24,400 = 50 + 4880 = 4930

exports = 6000

so net exports = 6000 - 4930

= 1070

since net exports are positive , there is a trade surplus

d) trade balance would occur when exports = imports

= 6000 = 50 + 0.2Y

5950 = 0.2Y

Y = 29,750

trade balance would occur when aggregate output = 29,750

(answered first 4 parts as per policy )